# Motives of nullcones of quiver representations

### v2

@@
Abstract.
@@
1. Introduction
For a finite dimensional complex representation of a complex reductive group, its nullcone, already introduced by Hilbert [6], plays an important role in the study of the invariant ring, the orbit structure, and the quotient map for the group action [7].
@@ 1. Introduction
In the present work, we study these nullcones quantitatively, by describing their motives, that is, their classes in the Grothendieck ring of complex varieties.
Our first result gives an explicit formula for the motives, as a recursion formula (Corollary 3.1), a generating function identity (Corollary 3.2), and as an explicit sum formula (Corollary 3.3). These formulas are derived from a geometric construction essentially using the stratification of nilpotent quiver representations by their socle type, inspired by methods of [12] (Section 2).
This result can be seen as a generalization of a prototypical motivic description of nullcones, namely the classical result [4] that there are $q^{d(d-1)}$ nilpotent $d\times d$-matrices over a finite field with $q$ elements.
Our second result (sections 4, 5) interprets the Hesselink stratification of [8] as an identity of motives, first yielding a positive summation formula (Theorem 5.2) for the motive of the nullcone, and culminating in a product formula relating the generating function of motives of nullcones to generating functions of semistable representations of level quivers (Theorem 5.7), reminiscent of wall-crossing formulas such as [11, Theorem 2.4].
It should be noted that the nullcone studied in this work is not to be confused with the nilpotent (Nakajima-type) quiver varieties studied quantitatively, in relation to Kac polynomials, in [1].
Acknowledgments: The second named author would like to thank S. Mozgovoy for stimulating discussions. The work of both authors is supported by the DFG GRK 2240 “Algebro-Geometric Methods in Algebra, Arithmetic and Topology”.
2. An explicit recursive formula for the motive of the nullcone
2.1. Quiver representations and their nullcone
[…]
$\langle{\bf d},{\bf e}\rangle=\sum_{i\in Q_{0}}d_{i}e_{i}-\sum_{\alpha:i\rightarrow j}d_{i}e_{j}$
@@ 2.1. Quiver representations and their nullcone
$R_{\mathbf{d}}=R_{\mathbf{d}}(Q)=\bigoplus_{\alpha:i\rightarrow j}{\rm Hom}_{\mathbb{C}}(V_{i},V_{j})$
@@
by $Q$-graded conjugation
$(g_{i})_{i}\cdot(V_{\alpha})_{\alpha}=(g_{j}V_{\alpha}g_{i}^{-1})_{\alpha:i\rightarrow j}.$
Let $N_{\mathbf{d}}$ be the nullcone for this action. By the its definition and the Hilbert criterion, it can be characterized by the following equivalent conditions:
Theorem 2.1.
[8, Section 1.4] The following are equivalent for a representation $V\in R_{\mathbf{d}}$:
@@
    all $G_{\bf d}$-invariants on $R_{\bf d}$ vanish at $V$, that is (by [10]), ${\rm tr}(V_{\alpha_{1}}\circ\ldots\circ V_{\alpha_{s}})=0$ for all oriented cycles $\alpha_{1}\ldots\alpha_{s}$ in $Q$,
@@
    all operators $V_{\alpha_{1}}\circ\ldots\circ V_{\alpha_{s}}$ for oriented cycles $\alpha_{1}\ldots\alpha_{s}$ in $Q$ are nilpotent,
@@
2.2. Motives
We denote by $K_{0}({\rm Var}_{\mathbb{C}})$ the Grothendieck ring of complex varieties; the class of a variety $X$ therein is denoted by $[X]$. We denote by $\mathbb{L}=[\mathbb{A}^{1}]$ the Lefschetz motive. We will work with the localized base ring of motives
$R=K_{0}({\rm Var}_{\mathbb{C}})[\mathbb{L}^{\pm 1},(1-\mathbb{L}^{i})^{-1}\,:\,i\geq 1].$
We define the Pochhammer symbol $(q)_{n}=(1-q)\cdot\ldots\cdot(1-q^{n})$. Using this notation, we have the following formulas for the motives of general linear groups and Grassmannians:
$[{\rm GL}_{n}(\mathbb{C})]=(-1)^{n}\mathbb{L}^{n(n-1)/2}(\mathbb{L})_{n},\;[{\rm Gr}_{k}^{n}]=\left[{n\atop k}\right]_{\mathbb{L}}:=\frac{(\mathbb{L})_{n}}{(\mathbb{L})_{k}\cdot(\mathbb{L})_{n-k}}.$
We also note that, for varieties $X$ and $Y$ admitting a bijective morphism $f:X\rightarrow Y$, we have $[X]=[Y]$ in $R$ ([2, Lemma 2.9]).
2.3. The recursive formula
Our aim is to describe various explicit formulas for the motive $[N_{\bf d}]\in R$. We introduce some shorthand notation concerning dimension vectors; namely, we define
@@ 2.3. The recursive formula
$\left[{{\bf d}\atop{{\bf d}^{1},\ldots,{\bf d}^{s}}}\right]_{q}=\prod_{i}\frac{(q)_{d_{i}}}{\prod_{k=1}^{s}(q)_{d^{k}_{i}}}$
for ${\bf d}={\bf d}^{1}+\ldots+{\bf d}^{s}\in\mathbb{N}Q_{0}$.
@@
We have the following recursion determining $[N_{\bf d}]$ for all ${\bf d}\not=0$:
@@
$\Phi_{i}(V):\bigoplus_{\alpha:i\rightarrow j}V_{\alpha}:V_{i}\rightarrow\bigoplus_{\alpha:i\rightarrow j}V_{j}.$
@@
$S_{\bf f}=\{(V_{\alpha})_{\alpha}\,|\,\dim{\rm Ker}(\Phi_{i}(V))=f_{i}\mbox{ for all }i\in Q_{0}\}.$
Since a nilpotent representation is an iterated extension of the one-dimensional simple representations $S_{i}$ of $Q$, it admits some $S_{i}$ as a subrepresentation, and thus belongs to some $S_{\bf f}$ for ${\bf f}\not={0}$. In other words, $V$ belongs to $S_{\bf f}$ if and only if the socle of $V$ is of dimension vector ${\bf f}$, which is nonzero since $V$ is so. In particular, we have $[S_{0}]=0$.
For ${\bf f}\leq{\bf d}$ we consider the product of Grassmannians
@@
and, more generally, for ${\bf f}\leq{\bf g}\leq{\bf d}$, the corresponding product ${\rm Fl}_{{\bf f},{\bf g}}({\bf d})$ of two-step flag varieties.
Inside ${\rm Gr}_{\bf f}({\bf d})\times N_{\bf d}$, we consider the closed subvariety
$X_{\bf f}=\{((U_{i})_{i},(V_{\alpha})_{\alpha})\,|\,U_{i}\subset{\rm Ker}(\Phi_{i}(V))\mbox{ for all }i\in Q_{0}\}.$
The first projection $p_{1}:X_{\bf f}\rightarrow{\rm Gr}_{\bf f}({\bf d})$ being equivariant for the obvious $G_{\bf d}$-actions, we have
$X_{\bf f}\simeq G_{\bf d}\times^{P_{\bf f}}Z_{\bf f},$
where, for a fixed choice of subspaces $(U_{i}^{0})_{i}\in{\rm Gr}_{\bf f}({\bf d})$, the group $P_{\bf f}\subset G_{\bf d}$ is the maximal parabolic fixing the $U_{i}^{0}$, and $Z_{\bf f}$ is the closed subset of $N_{\bf d}$ of representations for which every $V_{\alpha}$ for $\alpha:i\rightarrow j$ annihilates $U_{i}^{0}$. Choosing bases of the $V_{i}$ compatible with the $U_{i}^{0}$, in such representations the arrows of $Q$ are represented by block matrices
$\left[\begin{array}[]{cc}0&\zeta_{\alpha}\\ 0&W_{\alpha}\end{array}\right],$
where $(W_{\alpha})$ defines a representation in $N_{{\bf d}-{\bf f}}$. It follows that $Z_{\bf f}$ is a trivial vector bundle of rank
$\sum_{\alpha:i\rightarrow j}(d_{i}-f_{i})f_{j}$
over $N_{{\bf d}-{\bf f}}$. Consequently, we find for the motive of $X_{\bf f}$:
$[X_{\bf f}]=\frac{[G_{\bf d}]}{[P_{\bf f}]}\cdot[Z_{\bf f}]=\frac{[G_{\bf d}]}{[G_{{\bf d}-{\bf f}}]\cdot[G_{\bf f}]}\cdot\mathbb{L}^{-\langle{\bf d}-{\bf f},{\bf f}\rangle}\cdot[N_{{\bf d}-{\bf f}}].$
On the other hand, we consider the second projection $p_{2}:X_{\bf f}\rightarrow N_{\bf d}$. By definition of the stata $S_{\bf f}$, the image of $p_{2}$ is the union of the $S_{\bf g}$ for ${\bf f}\leq{\bf g}\leq{\bf d}$. We denote the inverse image of such a stratum $S_{\bf g}$ under $p_{2}$ by $X_{{\bf f},{\bf g}}$. By definition, the projection
$p_{2}:X_{{\bf f},{\bf g}}\rightarrow S_{\bf g}$
has all its fibres isomorphic to ${\rm Gr}_{\bf f}({\bf g})$. We claim that we have
$[X_{{\bf f},{\bf g}}]=[S_{\bf g}]\cdot[{\rm Gr}_{\bf f}({\bf g})]$
in $R$. For this, we observe that we have a diagram of Cartesian squares
$\begin{array}[]{ccccc}{\rm Fl}_{{\bf f},{\bf g}}({\bf d})&\leftarrow&\widetilde{X}&\stackrel{{\scriptstyle\widetilde{p}}}{{\rightarrow}}&X_{{\bf f},{\bf g}}\\ \downarrow&&\downarrow&&\downarrow\\ {\rm Gr}_{\bf g}({\bf d})&\leftarrow&X_{{\bf g},{\bf g}}&\stackrel{{\scriptstyle p}}{{\rightarrow}}&S_{\bf g}\end{array}$
@@
$\widetilde{X}=\{((U_{i})_{i},(U_{i}^{\prime})_{i},(V_{\alpha})_{\alpha})\,|\,U_{i}\subset U_{i}^{\prime}={\rm Ker}(\Phi_{i}(V))\mbox{ for all }i\in Q_{0}\}.$
As a special case of the previous observation, we see that the map $p:X_{{\bf g},{\bf g}}\rightarrow S_{\bf g}$ is a bijection, thus $\widetilde{p}$ is a bijection, too. The map ${\rm Fl}_{{\bf f},{\bf g}}({\bf d})\rightarrow{\rm Gr}_{\bf g}({\bf d})$ is Zariski-locally trivial due to $G_{\bf d}$-equivariance, thus the same holds for the map $\widetilde{X}\rightarrow X_{{\bf g},{\bf g}}$. It follows that we can calculate
$[X_{{\bf f},{\bf g}}]=[\widetilde{X}]=[X_{{\bf g},{\bf g}}]\cdot[{\rm Gr}_{\bf f}({\bf g})]=[S_{\bf g}]\cdot[{\rm Gr}_{\bf f}({\bf g})],$
@@
$[X_{\bf f}]=\sum_{{\bf f}\leq{\bf g}\leq{\bf d}}[S_{\bf g}]\cdot[{\rm Gr}_{\bf f}({\bf g})].$
We note the following identity of Gaussian binomial coefficients for all $m\geq 0$ (see for example [12, (3.14)]):
@@
$\sum_{{\bf f}\leq{\bf d}}(-1)^{\dim{\bf f}}\mathbb{L}^{\sum_{i}f_{i}(f_{i}-1)/2}[X_{\bf f}]=\sum_{{\bf f}\leq{\bf g}\leq{\bf d}}(-1)^{\dim{\bf f}}\mathbb{L}^{\sum_{i}f_{i}(f_{i}-1)/2}[{\rm Gr}_{\bf f}({\bf g})]\cdot[S_{\bf g}]=$
$=\sum_{{\bf g}\leq{\bf d}}\underbrace{(\sum_{{\bf f}\leq{\bf g}}(-1)^{\dim{\bf f}}\mathbb{L}^{\sum_{i}f_{i}(f_{i}-1)/2}[{\rm Gr}_{\bf f}({\bf g})])}_{=\delta_{{\bf g},{0}}}\cdot[S_{\bf g}]=[S_{0}]=0.$
Substituting the above formula for $[X_{\bf f}]$, we find
@@
Dividing by $[G_{\bf d}]$ and isolating the term corresponding to ${\bf f}={0}$, we find

### v3

@@
[…]
Motives of Nullcones of Quiver Representations
[…]
Motives of Nullcones of Quiver Representations
[…]
Received June 13, 2025, in final form February 10, 2026; Published online March 02, 2026
[…]
@@
[…]
nullcone; quiver representation; motive; Hesselink stratification
[…]
14C15; 14L24; 16G20
1 Introduction
For a finite-dimensional complex representation of a complex reductive group, its nullcone, already introduced by Hilbert [6], plays an important role in the study of the invariant ring, the orbit structure, and the quotient map for the group action [7].
@@ 1 Introduction
In the present work, we perform a systematic quantitative study of these nullcones, by describing their motives, that is, their classes in the Grothendieck ring of complex varieties. A first such formula (more precisely, a formula for counts of rational points over finite fields) already appears in [1, Lemma 2.6].
Our first result gives a simple explicit formula for the motives, as a recursion formula (Corollary 3.1), a generating function identity (Corollary 3.2), and as an explicit sum formula (Corollary 3.3). These formulas are derived from a geometric construction essentially using the stratification of nilpotent quiver representations by their socle type, inspired by methods of [12] (Section 2).
This result can be seen as a generalization of a prototypical motivic description of nullcones, namely the classical result [4] that there are $q^{d(d-1)}$ nilpotent $d\times d$ matrices over a finite field with $q$ elements.
Our second result (Sections 4 and 5) interprets the Hesselink stratification of [8] as an identity of motives, first yielding a positive summation formula (Theorem 5.2) for the motive of the nullcone, and culminating in a product formula relating the generating function of motives of nullcones to generating functions of semistable representations of level quivers (Theorem 5.7), reminiscent of wall-crossing formulas such as [11, Theorem 2.4].
It is reasonable to expect that our formulas also hold in positive characteristic, and thus yield formulas for numbers of rational points over finite fields. However, the methods of [2, 5] which we use in an essential way require a base field of characteristic zero.
2 An explicit recursive formula for the motive of the nullcone
2.1 Quiver representations and their nullcone
Let $Q$ be a finite quiver with set of vertices $Q_{0}$ and arrows written $\alpha\colon i\rightarrow j$. Define the Euler form $\langle\_,\_\rangle$ of $Q$ by
$\langle{\bf d},{\bf e}\rangle=\sum_{i\in Q_{0}}d_{i}e_{i}-\sum_{\alpha\colon i\rightarrow j}d_{i}e_{j}$
@@ 2.1 Quiver representations and their nullcone
$R_{\mathbf{d}}=R_{\mathbf{d}}(Q)=\bigoplus_{\alpha\colon i\rightarrow j}{\rm Hom}_{\mathbb{C}}(V_{i},V_{j})$
@@
by $Q$-graded conjugation $(g_{i})_{i}\cdot(V_{\alpha})_{\alpha}=\bigl(g_{j}V_{\alpha}g_{i}^{-1}\bigr)_{\alpha\colon i\rightarrow j}$.
Let $N_{\mathbf{d}}$ be the nullcone for this action. By its definition and the Hilbert criterion, it can be characterized by the following equivalent conditions.
Theorem 2.1 ([8, Section 1.4]).
The following are equivalent for a representation $V\in R_{\mathbf{d}}$:
@@ 2. $(2)$
    0 belongs to the closure of the orbit $G_{\bf d}\cdot V$,
@@ 3. $(3)$
    all $G_{\bf d}$-invariants on $R_{\bf d}$ vanish at $V$, that is ((by [10]), ${\rm tr}(V_{\alpha_{1}}\circ\dots\circ V_{\alpha_{s}})=0$ for all oriented cycles $\alpha_{1}\cdots\alpha_{s}$ in $Q$,
  4. $(4)$
    all operators $V_{\alpha_{1}}\circ\dots\circ V_{\alpha_{s}}$ for oriented cycles $\alpha_{1}\cdots\alpha_{s}$ in $Q$ are nilpotent,
@@ 5. $(5)$
2.2 Motives
We denote by $K_{0}({\rm Var}_{\mathbb{C}})$ the Grothendieck ring of complex varieties; the class of a variety $X$ therein is denoted by $[X]$. We denote by $\mathbb{L}=\big[\mathbb{A}^{1}\big]$ the Lefschetz motive. We will work with the localized base ring of motives
$\mathbb{M}=K_{0}({\rm Var}_{\mathbb{C}})\big[\mathbb{L}^{-1},\bigl(1-\mathbb{L}^{i}\bigr)^{-1}:i\geq 1\big]$
(the role of the localization is a purely formal one, to have the flexibility of dividing by motives of general linear groups). We define the Pochhammer symbol $(q)_{n}=(1-q)\cdots(1-q^{n})$. Using this notation, we have the following formulas for the motives of general linear groups and Grassmannians
${\rm GL}_{n}(\mathbb{C})]=(-1)^{n}\mathbb{L}^{n(n-1)/2}(\mathbb{L})_{n},\qquad[{\rm Gr}_{k}^{n}]=\left[{n\atop k}\right]_{\mathbb{L}}:=\frac{(\mathbb{L})_{n}}{(\mathbb{L})_{k}\cdot(\mathbb{L})_{n-k}}.$
We also note that, for varieties $X$ and $Y$ admitting a bijective morphism $f\colon X\rightarrow Y$, we have [X]=[Y]{[X]=[Y]} in $K_{0}({\rm Var}_{\mathbb{C}})$ [2, Lemma 2.9].
2.3 The recursive formula
Our aim is to describe various explicit formulas for the motive $N_{\bf d}]\in\mathbb{M}$. We introduce some shorthand notation concerning dimension vectors; namely, we define
@@ 2.3 The recursive formula
$\left[{{\bf d}\atop{{\bf d}^{1},\dots,{\bf d}^{s}}}\right]_{q}=\prod_{i}\frac{(q)_{d_{i}}}{\prod_{k=1}^{s}(q)_{d^{k}_{i}}}$
for ${\bf d}={\bf d}^{1}+\dots+{\bf d}^{s}\in\mathbb{N}Q_{0}$.
@@
We have the following recursion determining $N_{\bf d}]$ for all ${\bf d}\not=0$:
@@
$\Phi_{i}(V)=\bigoplus_{\alpha\colon i\rightarrow j}V_{\alpha}\colon\ V_{i}\rightarrow\bigoplus_{\alpha\colon i\rightarrow j}V_{j}.$
@@
$S_{\bf f}=\{(V_{\alpha})_{\alpha}\mid\dim{\rm Ker}(\Phi_{i}(V))=f_{i}\mbox{ for all }i\in Q_{0}\}.$
Since a nilpotent representation is an iterated extension of the one-dimensional simple representations $S_{i}$ of $Q$, it admits some $S_{i}$ as a subrepresentation, and thus belongs to some $S_{\bf f}$ for ${\bf f}\not={0}$}. In other words, $V$ belongs to $S_{\bf f}$ if and only if the socle of $V$ is of dimension vector $\bf f$}, which is nonzero since $V$ is so. In particular, we have $S_{0}]=0$.
For ${\bf f}\leq{\bf d}$, we consider the product of Grassmannians
@@
and, more generally, for ${\bf f}\leq{\bf g}\leq{\bf d}$, the corresponding product ${\rm Fl}_{{\bf f},{\bf g}}({\bf d})$ of two-step flag varieties. Inside ${\rm Gr}_{\bf f}({\bf d})\times N_{\bf d}$, we consider the closed subvariety
$X_{\bf f}=\{((U_{i})_{i},(V_{\alpha})_{\alpha})\mid U_{i}\subset{\rm Ker}(\Phi_{i}(V))\mbox{ for all }i\in Q_{0}\}.$
The first projection $p_{1}\colon X_{\bf f}\rightarrow{\rm Gr}_{\bf f}({\bf d})$ being equivariant for the obvious $G_{\bf d}$-actions, we have $X_{\bf f}\simeq G_{\bf d}\times^{P_{\bf f}}Z_{\bf f}$, where, for a fixed choice of subspaces $\bigl(U_{i}^{0}\bigr)_{i}\in{\rm Gr}_{\bf f}({\bf d})$, the group $P_{\bf f}\subset G_{\bf d}$ is the maximal parabolic fixing the $U_{i}^{0}$, and $Z_{\bf f}$ is the closed subset of $N_{\bf d}$ of representations for which every $V_{\alpha}$ for $\alpha\colon i\rightarrow j$ annihilates $U_{i}^{0}$. Choosing bases of the $V_{i}$ compatible with the $U_{i}^{0}$, in such representations the arrows of $Q$ are represented by block matrices
$\left[\begin{matrix}0&\zeta_{\alpha}\\ 0&W_{\alpha}\end{matrix}\right],$
where $(W_{\alpha})$ defines a representation in $N_{{\bf d}-{\bf f}}$. It follows that $Z_{\bf f}$ is a trivial vector bundle of rank $\sum_{\alpha\colon i\rightarrow j}(d_{i}-f_{i})f_{j}$ over $N_{{\bf d}-{\bf f}}$. Consequently, we find for the motive of $X_{\bf f}$
$X_{\bf f}]=\frac{[G_{\bf d}]}{[P_{\bf f}]}\cdot[Z_{\bf f}]=\frac{[G_{\bf d}]}{[G_{{\bf d}-{\bf f}}]\cdot[G_{\bf f}]}\cdot\mathbb{L}^{-\langle{\bf d}-{\bf f},{\bf f}\rangle}\cdot[N_{{\bf d}-{\bf f}}].$
On the other hand, we consider the second projection $p_{2}\colon X_{\bf f}\rightarrow N_{\bf d}$. By definition of the strata $S_{\bf f}$, the image of $p_{2}$ is the union of the $S_{\bf g}$ for ${\bf f}\leq{\bf g}\leq{\bf d}$. We denote the inverse image of such a stratum $S_{\bf g}$ under $p_{2}$ by $X_{{\bf f},{\bf g}}$. By definition, the projection $p_{2}\colon X_{{\bf f},{\bf g}}\rightarrow S_{\bf g}$ has all its fibres isomorphic to ${\rm Gr}_{\bf f}({\bf g})$. We claim that we have $X_{{\bf f},{\bf g}}]=[S_{\bf g}]\cdot[{\rm Gr}_{\bf f}({\bf g})]$ in $\mathbb{M}$. For this, we observe that we have a diagram of Cartesian squares
$\begin{array}[]{ccccc}{\rm Fl}_{{\bf f},{\bf g}}({\bf d})&\leftarrow&\widetilde{X}&\stackrel{{\scriptstyle\widetilde{p}}}{{\rightarrow}}&X_{{\bf f},{\bf g}}\\ \downarrow&&\downarrow&&\downarrow\\ {\rm Gr}_{\bf g}({\bf d})&\leftarrow&X_{{\bf g},{\bf g}}&\stackrel{{\scriptstyle p}}{{\rightarrow}}&S_{\bf g}.\end{array}$
@@
$\widetilde{X}=\{((U_{i})_{i},(U_{i}^{\prime})_{i},(V_{\alpha})_{\alpha})\mid U_{i}\subset U_{i}^{\prime}={\rm Ker}(\Phi_{i}(V))\mbox{ for all }i\in Q_{0}\}.$
As a special case of the previous observation, we see that the map $p\colon X_{{\bf g},{\bf g}}\rightarrow S_{\bf g}$ is a bijection (but we do not know a priori whether it is an isomorphism), thus $\widetilde{p}$ is a bijection, too. The map ${\rm Fl}_{{\bf f},{\bf g}}({\bf d})\rightarrow{\rm Gr}_{\bf g}({\bf d})$} is Zariski-locally trivial due to $G_{\bf d}$-equivariance, thus the same holds for the map $\widetilde{X}\rightarrow X_{{\bf g},{\bf g}}$. It follows that we can calculate
$X_{{\bf f},{\bf g}}]=\big[\widetilde{X}\big]=[X_{{\bf g},{\bf g}}]\cdot[{\rm Gr}_{\bf f}({\bf g})]=[S_{\bf g}]\cdot[{\rm Gr}_{\bf f}({\bf g})],$
@@
$X_{\bf f}]=\sum_{{\bf f}\leq{\bf g}\leq{\bf d}}[S_{\bf g}]\cdot[{\rm Gr}_{\bf f}({\bf g})].$
We note the following identity of Gaussian binomial coefficients for all $m\geq 0$ (see, for example, [12, equation (3.14)])
@@
$\sum_{{\bf f}\leq{\bf d}}(-1)^{\dim{\bf f}}\mathbb{L}^{\sum_{i}f_{i}(f_{i}-1)/2}[X_{\bf f}]$ $=\sum_{{\bf f}\leq{\bf g}\leq{\bf d}}(-1)^{\dim{\bf f}}\mathbb{L}^{\sum_{i}f_{i}(f_{i}-1)/2}[{\rm Gr}_{\bf f}({\bf g})]\cdot[S_{\bf g}]$
$=\sum_{{\bf g}\leq{\bf d}}\underbrace{\bigg(\sum_{{\bf f}\leq{\bf g}}(-1)^{\dim{\bf f}}\mathbb{L}^{\sum_{i}f_{i}(f_{i}-1)/2}[{\rm Gr}_{\bf f}({\bf g})]\bigg)}_{=\delta_{{\bf g},{0}}}\cdot[S_{\bf g}]$
$=[S_{0}]=0.$
Substituting the above formula for $X_{\bf f}]$, we find
@@
Dividing by $G_{\bf d}]$ and isolating the term corresponding to ${\bf f}={0}$, we find